can someone please show ma how to do this .thanks in advance for your help.find the roots of each quadratic equation by factoring.
1. x^2-6xi got 2,4
2.x^2-9/4x-25 i got 4x3-100x-9/4x

Answer by stanbon(75887)

find the roots of each quadratic equation by factoring.
1. x^2-6x =0 ; i got 2,4
x(x-6)=0
x = 0 or x=6
===========================
2.x^2-9/4x-25 = 0; i got 4x3-100x-9/4x
The fraction is zero when the numerator is zero.
x^2-9 = 0
(x-3)(x+3)=0
x = 3 or x = -3
================
Cheers,
Stan H.